OpenStudy (anonymous):

2C2H2 (g) + SO2 (g) → 4CO2 (g) +2H2O (l) Calculate the mass of CO2 produced from the complete combustion of 1.00g of C2H2. How do you do this?

OpenStudy (dls):

Is the answer 3.38 grams?

OpenStudy (anonymous):

YES! But how did you get there?

OpenStudy (anonymous):

You have some amazing knowledge you must pass on to me!

OpenStudy (chmvijay):

see here 2 moles of acetylene gives 4 mole of carbon dioxide mean 1 mole of aetylene gives 2 moles of carbon dioxide convert mole to their molecular weight that is 26.04 gram of acetylene gives 2 * 44.01 1 gram of acetylene gives =(2*44.01)/26.04=3.38

OpenStudy (chmvijay):

did u understand

OpenStudy (anonymous):

Thank you !! :D :D

OpenStudy (anonymous):

and yes :)

OpenStudy (dls):

Here is what I did: A simple approach. 2 moles of C2H2 give 4 moles of CO2 (g) 1 mole would give 2 moles Of CO2 right? Now, 1 mole of C2H2=12*2+2 = 26 grams 2 moles of CO2 = 44*2 = 88 grams so It means, 26 grams -----------give 88 grams 1 gram---------88/26=3.38 ! Hope you get it!

OpenStudy (chmvijay):

it is the same as i did sir @DLS

OpenStudy (dls):

ah yeah :P